Grade 12th passThermal Physics

at what temperature does the kelvin thermomrtr reads twice the fahrenheit thermometer

fariba
6 Years agoGrade 12th pass

2 Answers

Arun
6 Years ago

C = 5/9(F-32) or (F-32)/1.8

We want C = 2F or F = C/2

So C = (C/2 - 32)/1.8

1.8 C = C/2 - 32

1.3 C = -32

C = -24.6, F = -12.3

-24.6 degrees C = -12.3 degrees F. That’s rounded to the nearest tenth of a degree.

Aditya Gupta
6 Years ago
we know the relation F = 9/5(° K - 273) + 32
given that K= 2F
so F = 9/5(2F - 273) + 32
or F= 2297/13= 176.69 F or 353.53 K